OpenStudy (anonymous):

What is the fourth term of (d – 4b)^3? (1 point) b3 –b^3 64b^3 –64b^3

OpenStudy (anonymous):

(a-b)^3 = a^3 - 3a^2 b + 3 a b^2 - b^3

OpenStudy (anonymous):

i dont understand how im goin to use that to get an answer. im so lost with this question.

OpenStudy (anonymous):

U know binomial theorem, do you ?

OpenStudy (anonymous):

kind of but after i have (a-b)^3 = a^3 - 3a^2 b + 3 a b^2 - b^3 what do i do next?

OpenStudy (anonymous):

Here, what do u see as the fourth term .. -b^3 .. right?

OpenStudy (anonymous):

yes.